Find the currents and voltages in the circuit shown in Fig. 2.28.


1
Expert's answer
2022-04-04T09:05:47-0400

Assume

"R_1=R_2=1\Omega"

"R_{eq}=1+1=2\Omega"

V=2V



Current

"I=\frac{2}{2}=1A"

"I_1=I_2=1A"

Voltage

"V_1=I_1R_1=1\times1=1V"

"V_2=I_2R_2=1\times1=1V"